# A note on the Maxwell’s eigenvalues on thin sets

### v1

@@ 2.2. Hodge Laplacian and geometric functional setting
The weak formulation (2.2) and the functional setting described in Subsection 2.1, which seem apparently tied to Euclidean $3$-dimensional domains, can be translated in the context of general compact Riemannian manifolds.
@@
The metric $g$ allows to define a Hodge-star operator $\star:\Omega^{p}(M)\to\Omega^{n-p}(M)$. More concretely, we first note that the Riemannian metric induces a scalar product on the space of $p$-forms which we denote by $\langle\cdot,\cdot\rangle_{g}$. Then, for any $\omega\in\Omega^{p}(M)$, the Hodge-star operator $\star$ is defined by the following identity
@@
Hence the eigenvalue equation in (1.1) corresponds to $\Delta E=\lambda E$ when $E$ is co-closed, that is, when $\delta E=0$ (${\rm div}E=0$). Here, with abuse of notation, we have identified the vector field $E$ with its dual $1$-form. The boundary condition $\nu\times E=0$ on $\partial\Omega$ can be translated into $i^{*}E=0$ on $\partial\Omega$, where $i^{*}:\partial\Omega\to\Omega$ is the canonical inclusion. This condition forces $E$ to be normal to the boundary. In the terminology of differential forms, $i^{*}E=0$ on $\partial\Omega$ is called relative condition. In the case of 0-forms, the relative condition corresponds to Dirichlet boundary condition.
We now consider the functional setting for the Hodge Laplacian acting on forms. Having a scalar product induced by the metric on $\Omega^{p}(M)$, the definitions of $L^{2}$ spaces and Sobolev spaces extend naturally to differential $p$-forms: the space $L^{2}\Omega^{p}(M)$ is defined as the completion of $\Omega^{p}(M)$ with respect to the $L^{2}$ inner product on forms: $\int_{M}\langle\omega_{1},\omega_{2}\rangle_{g}dv_{g}$. The Sobolev spaces $H^{m}\Omega^{p}(M)$, $m\in\mathbb{N}$ are defined analogously, through the natural connection $\nabla$ on $(M,g)$ induced by the Riemannian metric (which allows to differentiate forms). It is then possible to define the analogous of the spaces $H_{0}({\rm curl},\Omega)$, $H({\rm div},\Omega)$, $H({\rm div}0,\Omega)$ on $(M,g)$ for differential forms of any degree. More precisely, we have
@@
We can see now that problem (2.2) corresponds to
@@
which holds for $\omega\in X_{N}(M,g)$. It is clear now that all the discussion in Subsection 2.1 applies to this more general setting, in fact the embedding $H^{1}\Omega^{p}(M)\to L^{2}\Omega^{p}(M)$ is compact. This means that problem (2.4) is associated with a compact, self-adjoint operator in $L^{2}\Omega^{p}(M)$ with non-negative, discrete spectrum. Gaffney’s inequality was originally proved in [24, 25] for manifolds without boundary, and in [23] for manifolds with boundary. For manifolds with boundary, it holds also under very mild smoothness assumptions on $\partial M$, see [33]. More precisely, for Euclidean domains, a Lipschitz condition on the boundary and a uniform outer ball condition are sufficient to guarantee the validity of Gaffney’s inequality.
For more details on Sobolev spaces of $p$-forms, Hodge-Morrey decomposition, Gaffney’s inequality, and for other relevant results of functional analysis, we refer to [42]. For more information on the specturm of the Hodge Laplacian on $p$-forms (with different type of boundary conditions) we refer e.g., to [27, 28, 37, 38].
The geometric framework described above allows for a more general approach to eigenvalue problems of the type (1.1) and helps avoiding some technicalities that may arise from the explicit use of coordinates. It gives a geometric meaning to the decomposition of the Maxwell spectrum in the ${\rm TE},{\rm TM}$ and $\rm TEM$} modes for domains $\omega\times I$ (where $\omega\subset\mathbb{R}^{2}$ and $I\subset\mathbb{R}$) contained in [15]. Moreover, even though it is not the purpose of the present paper, it can be applied in any dimension and any ambient Riemannian manifold.
We refer to [27] for a nice introduction to eigenvalue problems for the Laplacian on $p$ forms on manifolds with boundary. See also [40] for a detailed exposition on the spectrum of the Laplacian on $p$-forms on convex Euclidean domains.
@@
Throughout this section we shall denote by $M$ the following product manifold of dimension $3$:
@@
where $(\Sigma,g_{\Sigma})$ is a compact Riemannian surface and $I=(0,h)$ is an interval, $h>0$. The surface $\Sigma$ is compact, connected, smooth, and is allowed to have a smooth boundary (with any number of connected components) and a possibly non-trivial topology. By $g_{\Sigma}$ we denote a smooth Riemannian metric on $\Sigma$. By $x$ we shall denote a point of $\Sigma$ and by $t\in(0,h)$ the usual coordinate on $I$.
@@
restricted to the space of co-closed $1$-forms, that is, $1$-forms $\omega$ verifying $\delta\omega=0$ . To be more precise, we should have written $\delta_{g_{p}}$ and $\Delta_{g_{p}}$ since the co-differential depends on the metric. We shall drop the subscript when it is clear from the context. Here $i:\partial M\to M$ is the canonical inclusion. This problem is the eigenvalue problem for the Hodge Laplacian restricted to co-closed one-forms with relative boundary conditions, see [27].
We describe now the eigenvalues and eigenfunctions of (3.1) on the product manifold $M$. We start from the case $\partial\Sigma=\emptyset$.
@@ 3.1. Eigenvalues and eigenfunctions when $\partial\Sigma=\emptyset$
    $F_{jk}(x,t)=H_{k}(x)u_{j}(t)$, where $\{H_{k}\}_{k=1}^{2g}$ is a basis of harmonic $1$-forms on $\Sigma$.
@@
We can recognize the three families of modes described in [15] for cylinders and balls in $\mathbb{R}^{3}$. In particular, our first family of corresponds to the $\rm TE$} modes in [15], the second family corresponds to the $\rm TM$} modes, while, if the topology is not trivial, our third family corresponds to the $\rm TEM$} modes.
@@
Here and in what follow, by $\Delta_{\Sigma}$ we denote the Laplacian on $\Sigma$ for the metric $g_{\Sigma}$ and by $\delta_{\Sigma}$ we denote the codifferential on $\Sigma$ for the metric $g_{\Sigma}$. Therefore we need to solve
@@
(3.3) $\begin{cases}h\Delta^{2}_{\Sigma}w-h^{\prime\prime}\Delta_{\Sigma}w-\lambda h\Delta_{\Sigma}w=0\,,&{\rm in\ }M\\ v^{\prime}(0)d_{\Sigma}w=v^{\prime}(h)d_{\Sigma}w=0\,,&{\rm on\ }\Sigma\times\{0,h\}.\end{cases}$
If $w={\rm const}$ on $\Sigma$ we would have $F=0$. Moreover we can choose $w$ such that $\int_{\Sigma}w=0$, since the corresponding $F$ would not change. Hence the boundary condition reads $v^{\prime}(0)=v^{\prime}(h)=0$.
@@
$\begin{cases}-u^{\prime\prime}d_{\Sigma}w+hd_{\Sigma}\Delta_{\Sigma}w=\lambda ud_{\Sigma}w\,,&{\rm in\ }M\\ u(0)d_{\Sigma}w=u(h)d_{\Sigma}w=0.\end{cases}$
@@ 3.2. Eigenvalues and eigenfunctions when $\partial\Sigma\neq\emptyset$
Summarizing, for any $u\in V_{F}$, there exists a unique $\tilde{u}\in V_{p}$, $\tilde{u}=u+dv$, with $\delta_{g_{p}}\tilde{u}=0$, $du=d\tilde{u}$, and $\tilde{u}$ is tangential for $g_{p}$. This last fact follows since for a solution $v$ of (4.3), $dv$ is tangential, and since a $1$-form $u$ is tangential for $g_{F}$ if and only if it is tangential for $g_{p}$. This is due to the structure of the metrics $g_{F},g_{p}$ which, in local coordinates, assume the form of a block diagonal matrix splitting the components along $T\Sigma$ and $(0,h)$. Hence we can write
@@ 5. Convergence of eigenfunctions
Let $\mu_{j}$ be an eigenvalue of the Laplacian on $\Sigma$ (with Dirichlet conditions if $\partial\Sigma\neq\emptyset$) and let $w$ be the associated eigenfunction. Then there exists $h_{0}>0$ and an eigenfunction $u$ of problem (5.1) on $(M,g_{F})$ associated with $\Lambda_{j}(h,g_{F})$ satisfying
@@
Let $\mu_{j}$ be an eigenvalue of the Laplacian on $\Sigma$ (with Dirichlet conditions if $\partial\Sigma\neq\emptyset$) and let $w$ be the associated eigenfunction. Then there exists $h_{0}>0$ and an eigenfunction $u$ of problem (3.1) associated with $\lambda_{j}(h,g_{F})$ satisfying
@@
This last corollary, translated in terms of solutions of Maxwell problem (1.1), says that, for all $h>0$ sufficiently small, given an eigenfunction $w$ on the limit surface $\Sigma$ associated with a limit eigenvalue $\mu_{j}$ (or $\mu_{j}^{D}$ if $\Sigma$ has a boundary), there exists an eigenfunction $u$ of (1.1) associated with $\lambda_{j}(\Omega_{h})$ which is close in $L^{2}(\Omega)^{3}$ to the constant extension of $w$ in the normal direction to $\Sigma$.
@@ Appendix A Full Hodge Laplacian spectrum with relative conditions on product manifolds
We have seen that the Maxwell eigenvalues coincide with the eigenvalues of the Hodge Laplacian with relative conditions restricted to the subspace of co-closed differential forms. On a compact Riemannian manifold $(M,g)$, the eigenvalue problem for the Hodge Laplacian with relative conditions acting on $p$ forms reads
@@
Note that the Hodge Laplacian spectrum restricted to co-closed $1$-forms (Maxwell spectrum) is a subset of the Hodge Laplacian spectrum. However, it is not immediate to recognize the explicit form of the co-closed eigenfunctions as in Theorems 3.1 and 3.4. Nevertheless, if we denote by $\Lambda_{j}(h,g_{p})$ the eigenvalues of the (full) Hodge Laplacian on $(M,g_{p})$ with relative conditions, then, for all $j\in\mathbb{N}$,
@@ Acknowledgments
The authors are very thankful to Prof. Alessandro Savo for the numerous discussions on the topic and for reading a preliminary version of the paper, providing helpful advice.

### v2

@@ 2.2. Hodge Laplacian and geometric functional setting
The weak formulation (2.2) and the functional setting described in Subsection 2.1, which apparently seem tied to Euclidean $3$-dimensional domains, can be translated in the context of general compact Riemannian manifolds.
@@
The metric $g$ allows us to define a Hodge-star operator $\star:\Omega^{p}(M)\to\Omega^{n-p}(M)$. More concretely, we first note that the Riemannian metric induces a scalar product on the space of $p$-forms which we denote by $\langle\cdot,\cdot\rangle_{g}$. Then, for any $\omega\in\Omega^{p}(M)$, the Hodge-star operator $\star$ is defined by the following identity
@@
Hence, the eigenvalue equation in (1.1) corresponds to $\Delta E=\lambda E$ when $E$ is co-closed, that is, when $\delta E=0$ (${\rm div}E=0$). Here, with abuse of notation, we have identified the vector field $E$ with its dual $1$-form. The boundary condition $\nu\times E=0$ on $\partial\Omega$ can be translated into $i^{*}E=0$ on $\partial\Omega$, where $i^{*}:\partial\Omega\to\Omega$ is the canonical inclusion. This condition forces $E$ to be normal to the boundary. In the terminology of differential forms, $i^{*}E=0$ on $\partial\Omega$ is called relative condition. In the case of 0-forms, the relative condition corresponds to Dirichlet boundary condition.
We now consider the functional setting for the Hodge Laplacian acting on forms. Having a scalar product induced by the metric on $\Omega^{p}(M)$, the definitions of $L^{2}$ spaces and Sobolev spaces extend naturally to differential $p$-forms: the space $L^{2}\Omega^{p}(M)$ is defined as the completion of $\Omega^{p}(M)$ with respect to the $L^{2}$-inner product on forms: $\int_{M}\langle\omega_{1},\omega_{2}\rangle_{g}dv_{g}$. The Sobolev spaces $H^{m}\Omega^{p}(M)$, $m\in\mathbb{N}$ are defined analogously, through the natural connection $\nabla$ on $(M,g)$ induced by the Riemannian metric (which allows to differentiate forms). It is then possible to define the analogous of the spaces $H_{0}({\rm curl},\Omega)$, $H({\rm div},\Omega)$, $H({\rm div}0,\Omega)$ on $(M,g)$ for differential forms of any degree. More precisely, we have
@@
We can now see that problem (2.2) corresponds to
@@
which holds for $\omega\in X_{N}(M,g)$. It is clear now that all the discussion in Subsection 2.1 applies to this more general setting, since the embedding $H^{1}\Omega^{p}(M)\to L^{2}\Omega^{p}(M)$ is compact. This means that problem (2.4) is associated with a compact, self-adjoint operator in $L^{2}\Omega^{p}(M)$ with nonnegative, discrete spectrum. Gaffney’s inequality was originally proved in [24, 25] for manifolds without boundary, and in [23] for manifolds with boundary. For manifolds with boundary, it holds also under very mild smoothness assumptions on $\partial M$, see [33]. More precisely, for Euclidean domains, a Lipschitz condition on the boundary and a uniform outer ball condition are sufficient to guarantee the validity of Gaffney’s inequality.
For more details on Sobolev spaces of $p$-forms, Hodge-Morrey decomposition, Gaffney’s inequality, and for other relevant results of functional analysis, we refer to [42]. For more information on the spectrum of the Hodge Laplacian on $p$-forms (with different types of boundary conditions) we refer, e.g., to [27, 28, 37, 38].
The geometric framework described above allows for a more general approach to eigenvalue problems of type (1.1) and helps to avoid some technicalities that may arise from the explicit use of coordinates. It gives a geometric meaning to the decomposition of the Maxwell’s spectrum in the ${\rm TE},{\rm TM}$ and $\rm TEM$} modes for domains $\omega\times I$ (where $\omega\subset\mathbb{R}^{2}$ and $I\subset\mathbb{R}$) contained in [15]. Moreover, even though it is not the purpose of the present paper, it can be applied in any dimension and any ambient Riemannian manifold.
We refer to [27] for a nice introduction to eigenvalue problems for the Laplacian on $p$-forms on manifolds with boundary. See also [40] for a detailed exposition on the spectrum of the Laplacian on $p$-forms on convex Euclidean domains.
@@
Throughout this section, we shall denote by $M$ the following product manifold of dimension $3$:
@@
where $(\Sigma,g_{\Sigma})$ is a compact Riemannian surface, $I=(0,h)$ is an interval, and $h>0$. The surface $\Sigma$ is compact, connected, smooth, and it is allowed to have a smooth boundary (with any number of connected components) and a possibly non-trivial topology. By $g_{\Sigma}$ we denote a smooth Riemannian metric on $\Sigma$. By $x$ we shall denote a point of $\Sigma$ and by $t\in(0,h)$ the usual coordinate on $I$.
@@
restricted to the space of co-closed $1$-forms, that is, $1$-forms $\omega$ verifying $\delta\omega=0$ . To be more precise, we should have written $\delta_{g_{p}}$ and $\Delta_{g_{p}}$ since the co-differential depends on the metric. We shall omit the subscript when its meaning is clear from the context. Here $i:\partial M\to M$ is the canonical inclusion. This problem is the eigenvalue problem for the Hodge Laplacian restricted to co-closed one-forms with relative boundary conditions, see [27].
We now describe the eigenvalues and eigenfunctions of (3.1) in the product manifold $M$. We start from the case $\partial\Sigma=\emptyset$.
@@ 3.1. Eigenvalues and eigenfunctions when $\partial\Sigma=\emptyset$
    $F_{jk}(x,t)=H_{k}(x)u_{j}(t)$, where $\{H_{k}\}_{k=1}^{2\gamma}$ is a basis of harmonic $1$-forms on $\Sigma$.
@@
We can recognise the three families of modes described in [15] for cylinders and balls in $\mathbb{R}^{3}$. In particular, our first family of corresponds to the $\rm TM$} modes in [15], the second family corresponds to the $\rm TE$} modes, while, if the topology is not trivial, our third family corresponds to the $\rm TEM$} modes.
@@
Here and in what follows, by $\Delta_{\Sigma}$ we denote the Laplacian on $\Sigma$ for the metric $g_{\Sigma}$ and by $\delta_{\Sigma}$ we denote the codifferential on $\Sigma$ for the metric $g_{\Sigma}$. Therefore, we need to solve
@@
(3.3) $\begin{cases}v\Delta^{2}_{\Sigma}w-v^{\prime\prime}\Delta_{\Sigma}w-\lambda v\Delta_{\Sigma}w=0\,,&{\rm in\ }M\\ v^{\prime}(0)d_{\Sigma}w=v^{\prime}(h)d_{\Sigma}w=0\,,&{\rm on\ }\Sigma\times\{0,h\}.\end{cases}$
If $w={\rm const}$ on $\Sigma$ we would have $F=0$. Moreover, we can choose $w$ such that $\int_{\Sigma}w=0$, since the corresponding $F$ would not change. Hence, the boundary condition reads $v^{\prime}(0)=v^{\prime}(h)=0$.
@@
$\begin{cases}-u^{\prime\prime}d_{\Sigma}w+u\,d_{\Sigma}\Delta_{\Sigma}w=\lambda ud_{\Sigma}w\,,&{\rm in\ }M\\ u(0)d_{\Sigma}w=u(h)d_{\Sigma}w=0.\end{cases}$
@@ 3.2. Eigenvalues and eigenfunctions when $\partial\Sigma\neq\emptyset$
Summarizing, for any $u\in V_{p}$, there exists a unique $\tilde{u}\in V_{F}$, $\tilde{u}=u+dv$, with $\delta_{g_{F}}\tilde{u}=0$, $du=d\tilde{u}$, and $\tilde{u}$ is tangential for $g_{F}$. This last fact follows since for a solution $v$ of (4.3), $dv$ is tangential, and since a $1$-form $u$ is tangential for $g_{F}$ if and only if it is tangential for $g_{p}$. This is due to the structure of the metrics $g_{F},g_{p}$ which, in local coordinates, assume the form of a block diagonal matrix splitting the components along $T\Sigma$ and $(0,h)$. Hence we can write
@@ 5. Convergence of eigenfunctions
Let $\mu_{j}$ be an eigenvalue of the Laplacian on $\Sigma$ (with Dirichlet conditions if $\partial\Sigma\neq\emptyset$) and let $w$ be an associated $L^{2}$-normalized eigenfunction. Then there exists $h_{0}>0$ and an eigenfunction $u$ of problem (5.1) on $(M,g_{F})$ associated with $\Lambda_{j}(h,g_{F})$ satisfying
@@
Let $\mu_{j}$ be an eigenvalue of the Laplacian on $\Sigma$ (with Dirichlet conditions if $\partial\Sigma\neq\emptyset$) and let $w$ be an associated $L^{2}$-normalized eigenfunction. Then there exists $h_{0}>0$ and an eigenfunction $u$ of problem (3.1) associated with $\lambda_{j}(h,g_{F})$ satisfying
@@
This last corollary, translated in terms of solutions of Maxwell’s problem (1.1), says that, for all $h>0$ sufficiently small, given an eigenfunction $w$ on the limit surface $\Sigma$ associated with a limit eigenvalue $\mu_{j}$ (or $\mu_{j}^{D}$ if $\Sigma$ has a boundary), there exists an eigenfunction $u$ of (1.1) associated with $\lambda_{j}(\Omega_{h})$ which is close in $L^{2}(\Omega)^{3}$ to the constant extension of $w$ in the normal direction to $\Sigma$.
@@
Remark 5.3.
Concerning stronger notions of convergence, it is natural to expect that the resolvent operator of the Maxwell’s system converges to that of the (Dirichlet) Laplacian on $\Sigma$. We believe that this is the case. However, the proof of this fact is beyond the scope of the paper which is devoted to the analysis of the eigenvalues and the consequences for the spectral geometry of this operator. Although the convergence in the case of the product metric could be achieved with standard arguments, the convergence in the Euclidean setting is more delicate. Indeed, the natural extension operator, taking a function $w$ on $\Sigma$ to a $1$-form wdtwdt in $M$, does not produce, in general, co-closed forms; and even more dramatically, the projection of wdtwdt on the space of co-closed forms may depend on the small parameter $h$.
@@ Appendix A Full Hodge Laplacian spectrum with relative conditions on product manifolds
We have seen that the Maxwell’s eigenvalues coincide with the eigenvalues of the Hodge Laplacian with relative conditions restricted to the subspace of co-closed differential forms. On a compact Riemannian manifold $(M,g)$, the eigenvalue problem for the Hodge Laplacian with relative conditions acting on $p$ forms reads
@@
Note that the Hodge Laplacian spectrum restricted to co-closed $1$-forms (Maxwell’s spectrum) is a subset of the Hodge Laplacian spectrum. However, it is not immediate to recognize the explicit form of the co-closed eigenfunctions as in Theorems 3.1 and 3.4. Nevertheless, if we denote by $\Lambda_{j}(h,g_{p})$ the eigenvalues of the (full) Hodge Laplacian on $(M,g_{p})$ with relative conditions, then, for all $j\in\mathbb{N}$,
@@ Acknowledgments
The authors are very thankful to Prof. Alessandro Savo for the numerous discussions on the topic and for reading a preliminary version of the paper, providing helpful advice, and to Prof. Francesco Bei for his helpful advice.